A certain mass of a gass occupies 39ml at 760mm pressures.What volume would it occupies if the pressure is raised to 760ml provided that remains constant?
 

Grade:11

1 Answers

Arun
6 years ago
Dear Raja
 
If temp is consrant then
 
P1 * V1 = P2 * V2
760 * 39 = 780 * V2
V2 = 38 ml
 
Regards
Arun (askIITians forum expert)